### v2

Abstract
Turbulent-flow control aims to develop strategies that effectively manipulate fluid systems, such as the reduction of drag in transportation and enhancing energy efficiency, both critical steps towards reducing global CO2 emissions. Deep reinforcement learning (DRL) offers novel tools to discover flow-control strategies, which we combine with our knowledge of the physics of turbulence. We integrate explainable deep learning (XDL) to objectively identify the coherent structures containing the most informative regions in the flow, with a DRL model trained to reduce them. The trained model targets the most relevant regions in the flow to sustain turbulence and produces a drag reduction which is higher than that of a model specifically trained to reduce the drag, while using only half its power consumption. Moreover, the XDL model results in a better drag reduction than other models focusing on specific classically identified coherent structures. This demonstrates that combining DRL with XDL can produce causal control strategies that precisely target the most influential features of turbulence. By directly addressing the core mechanisms that sustain turbulence, our approach offers a powerful pathway towards its efficient control, which is a long-standing challenge in physics with profound implications for energy systems, climate modeling and aerodynamics.
Turbulence remains one of the last unsolved problems of classical physics [1], defying a complete theoretical description despite its ubiquitous presence in natural [2] and engineering [3] flows. In fact, turbulent fluid flows are present in a myriad of industrial applications: from the food industry to the cooling of microprocessors, or harnessing wind power. Turbulence is also the primary cause of viscous drag with $30\%$ of the energy consumption worldwide being spent on overcoming drag in transportation [4]. Thus, improving the aerodynamic efficiency of terrestrial and airborne vehicles plays a pivotal role in the global reduction of CO2 emissions. Despite centuries of theoretical [5], experimental [6], and computational [7] progress since Leonardo da Vinci’s first observations [8], many fundamental questions on the nature of turbulence remain unsolved [9, 10, 11], preventing us from efficiently manipulating turbulent flows.
Figure 1: The overall architecture of the deep reinforcement learning framework. The scheme shows the communication between the involved parts. The top shows the information inputs to the DRL agent, i.e. the rewards ($r_{t}$) and states ($s_{t}$) at each time $t$, and outputs to the flow, i.e. the actions ($a_{t}$). In this multi-agent reinforcement-learning (MARL) framework, the agents cooperate featuring a shared neural network. The middle panel depicts the computational fluid dynamics (CFD) environment, from which each agent receives the information $(u,v)$ of the point above within a plane located at $y^{+}=15$. The actions are applied at the wall in the form of blowing and suction. At the bottom, we illustrate the different rewards considered. We note that the SHAP values are computed instantaneously from the velocity field through a pre-trained U-net model.
In this work, we incorporate the most recent progress in understanding the underlying dynamics of turbulence to design novel control strategies which reduce the friction produced by the flow. In the chaotic motion that characterizes wall-bounded turbulence, it is possible to identify persistent spatio-temporal patterns (coherent structures) that interact in the near-wall cycle of turbulence [12, 13]. This cycle involves streamwise vortices generating elongated velocity perturbations (streaks), which subsequently experience instabilities and breakdown, generating new vortices and giving rise to a self-sustaining process (SSP) [14]. Existing approaches to drag reduction, such as opposition control [15], seek to disrupt this SSP based on heuristic observations, for instance by targeting ejection and sweep events [16]. Recently, the advent of new data-intensive techniques has expanded the available toolbox, with deep reinforcement learning (DRL) emerging as a particularly powerful approach to reveal new, more effective control strategies. DRL has demonstrated success across diverse physical domains, as broad as plasma fusion [17], optics [18], and fluid dynamics. Within this domain, DRL has been successfully applied to a wide range of control problems, including coordinated motion in fish schooling [19] and the control of separated flows [20, 21].
The framework for DRL control is illustrated in Fig. 1: the DRL agent (or controller) receives partial observations $s_{t}$ of the system and is trained to generate actions $a_{t}$, which modify the turbulent flow in the effort to maximize the reward $r_{t}$. While previous DRL applications to turbulence control have targeted directly the reduction of turbulent drag [22, 23], our approach targets the SSP of near-wall turbulence, with rewards based on the agent’s ability to manipulate turbulent coherent structures.
We identify these structures using both conventional criteria (c.f. [16] and references therein) and data-driven methods based on explainable deep learning (XDL) [24]. Conventional coherent structures in turbulence, i.e. Q events, streaks, and vortices, are characterized using instantenous velocity fluctuations and their derivatives, while data-driven structures are identified using Shapley additive explanation (SHAP) values [25], which are obtained through a game-theoretic method that calculates the importance of each input feature in a deep neural network. When applied to a U-net based on convolutional layers used for prediction of the future states of the flow, the SHAP values successfully identify regions of importance in wall-bounded turbulence [24].
The present study demonstrates how XDL can effectively be coupled with DRL to improve the performance of the learned policies compared with conventional physics-based rewards. We evaluate four reward strategies: (i) wall-shear stress reduction, (ii) Q-event reduction, (iii) streak reduction, and (iv) reduction of the SHAP values. The first reward is the most direct approach, rewards (ii) and (iii) target coherent structures traditionally studied in the turbulence literature, while (iv) is entirely data-driven yet physics-aware via XDL. To design the rewards, we choose appropriate instantaneous proxies to characterize the intensity of the turbulent coherent structures. In the SHAP-based reward, a U-net is used to predict the SHAP values from the instantaneous velocity field, as discussed in more detail in the Methods section.
Our work presents an entirely data-driven method that requires no prior flow hypotheses, and yet is capable of identifying the physical processes governing wall-bounded turbulence dynamics. The implications of the present study extend beyond turbulence control, suggesting a powerful tool where XDL and DRL can uncover fundamental physical mechanisms that might otherwise remain out of reach using traditional approaches.
Targeting coherent structures for drag reduction
Figure 2: Reduction of the quantities of interest with respect to the uncontrolled case. (Left) Drag reduction, (top-right) proxy for Q-event reduction and (bottom-right) proxy for streak reduction. Results are averaged over the 50 initial conditions used for policy evaluation. Solid lines denote mean values and shaded regions indicate one standard deviation. Colors indicate different control strategies: opposition control (blue), DRL for direct drag reduction (orange), DRL for Q-event reduction (green), DRL for streak reduction (red), and DRL for SHAP reduction (purple).
Our DRL set-up strategically identifies coherent structures to effectively influence the SSP of near-wall turbulence. This approach requires quantitative proxies to characterize the coherent structures in the flow. We define the instantaneous velocity vector $\mathbf{u}(x,y,z,t)=(u,v,w)$ along the streamwise, wall-normal and spanwise directions, where $t$ denotes time. Deviations from the instantaneous spatial mean averaged in the periodic directions are represented by primed variables: $u^{\prime}(x,y,z,t)=u(x,y,z,t)-\iint u(x,y,z,t)\mathrm{d}x\mathrm{d}z/(L_{x}L_{z})$ and $L_{x}$ and $L_{z}$ denote the domain sizes. We consider a turbulent open channel which can be characterized by the friction Reynolds number $Re_{\tau}=u_{\tau}h/\nu$, where $\nu$ denotes the fluid kinematic viscosity, $h$ the channel height, and $u_{\tau}=\sqrt{\tau_{w}/\rho}$ is the friction velocity (defined in terms of the wall-shear stress $\tau_{w}$ and the fluid density $\rho$). We consider $Re_{\tau}=180$ in our uncontrolled simulations. Quantities non-dimensionalized with the viscous scales $u_{\tau}$ and $\nu$ are denoted by the superscript ``+"``"``+"` ` + ". The DRL-discovered control strategies will be compared with the heuristic opposition control. Opposition control is a closed-loop control, where the control action is based on real-time flow-state sensing [15], which reduces friction drag by using blowing and suction with a vertical velocity distribution defined as:
$v_{\mathrm{wall}}(x,z,t)=-v^{\prime}(x,y_{s},z,t).$ (1)
This control opposes the vertical velocity on a plane at a wall-normal distance $y_{s}$, with the aim of suppressing the streamwise vortices.
The following control inputs are sampled at $y^{+}=15$: $v^{\prime}(t,x,y^{+}=15,z)$ for opposition control and $\{u^{\prime}(t,x,y^{+}=15,z),v^{\prime}(t,x,y^{+}=15,z)\}$ for the DRL-based techniques. To quantify the total intensity of Q events we consider the Reynolds stress averaged over the domain volume:
$\int_{\Omega}|(u(t,x,y,z)-U_{T}(y))v(t,x,y,z)|\rm{d}\Omega,$ (2)
where $U_{T}$ is the long-time spatio-temporal average for the turbulent uncontrolled case and $\Omega$ denotes the volume. Following [26], we characterize the streak intensity as:
$\int_{\Omega}\sqrt{u^{\prime}(t,x,y,z)^{2}+w^{\prime}(t,x,y,z)^{2}}\rm{d}\Omega.$ (3)
Similarly, we define the presence of highly informative regions based on the SHAP values as:
$\int_{\Omega}\sqrt{\phi_{x}^{\prime}(t,x,y,z)^{2}+\phi_{y}^{\prime}(t,x,y,z)^{2}+\phi_{z}^{\prime}(t,x,y,z)^{2}}\rm{d}\Omega,$ (4)
where $\bm{\phi}=(\phi_{x},~{}\phi_{y},~{}\phi_{z})$ represent the SHAP vector field as detailed in the Methods section. Here, the SHAP values are inferred instantaneously from the velocity field by a U-net trained on the uncontrolled channel data (see the Methods section for more details).
We consider two channel domain sizes in our study, one for training and one for testing. The DRL agents are trained in a small channel configuration (SCC), the so-called minimal flow unit [27], which is computationally affordable and supports a single SSP. The learnt policy is then tested in a large channel configuration (LCC) where many SSPs are present and interact. Here we discuss the performance in the larger channel, while the details of the computational settings and the SCC results are reported in the Methods section and in the Supplementary material, respectively.
The discovered policies exhibit significant differences between the various proposed rewards. Fig. 2 shows various performance metrics averaged over 50 different initial conditions. Our results exhibit a significant drag reduction: approximately $31.7\%$ when using DRL aiming specifically at direct drag (DD) reduction, $27.5\%$ when targeting Q events, and $21.2\%$ when targeting streaks. However, all these approaches show a worse performance than the one of the SHAP-based policy, which presents a $33.3\%$ drag reduction: $5\%$ more (1.6 percentage points) than DD-based control, $20.8\%$ more (5.6 percentage points) than Q-event-based control, and $56.8\%$ more (12.1 percentage points) than streak-based control. This finding is particularly remarkable given that all policies were initially trained in a small channel where performance differences were small (see the Methods section and Supplementary material).
We obtain further insights into the dynamic effect of the DRL-discovered control strategies when comparing the averaged Reynolds stresses and streak intensity in the various controlled cases, as shown in Fig. 2 (right). Note that the DRL approach targeting Q events exhibits the best performance reducing both Q events and streaks. Interestingly, the SHAP-based policy yields the highest drag reduction, despite not being particularly effective at reducing Q events or streaks, highlighting the fact that these classically studied structures are not the most important to target when reducing drag. The effectiveness of the SHAP-based policy stems from its ability to identify the most informative flow regions to predict future turbulence fluctuations in an objective manner. The DRL approach targeting SHAP values manipulates the most relevant structures for the flow evolution and the SSP of near-wall turbulence. By targeting these regions, the DRL control learns to hamper turbulence-supporting mechanisms in a causally informed manner, unlike the more simplistic approach of targeting the classically-studied coherent structures. These findings confirm that the traditional coherent structures provide an incomplete picture of the mechanisms supporting turbulence, and show that XDL complements and expands our understanding of wall-bounded turbulence [28].
Figure 3: Joint probability density function of the streamwise and wall-normal fluctuations at $y^{+}=15$ for the different control strategies. Results are averaged over the 50 initial conditions used for policy evaluation for: uncontrolled flow (black), opposition control (blue), DRL for DD (orange), DRL for Q events (green), DRL for streaks (red), and DRL for SHAP values (purple). Grey lines in all the panels show the uncontrolled case for comparison.
Fig. 3 shows the joint probability density function of streamwise ($u^{\prime}$) and wall-normal ($v^{\prime}$) perturbations on the sensing plane $y^{+}=15$ evaluated on 50 different trajectories. Note that the predominance of ejections and sweeps in the uncontrolled case is consistent with the wall-bounded turbulence literature [29] and the results from opposition control and DRL for drag reduction agree with our previous results [22]. Furthermore, the approaches based on classical structures produce distinctly different fluctuations intensities: Q-event-based control substantially reduces the perturbation intensities in both stream- and spanwise directions, while the streak-based control decreases streamwise perturbations but increases wall-normal perturbations, emphasizing the negative fluctuations. In contrast, our SHAP-based control leads to a quite even distribution of the four quadrants, with a slight preference towards positive streamwise fluctuations.
Figure 4: Power input and time signal of a single controller for the various control strategies. (Top left) Power input averaged over the 50 initial conditions used for policy evaluation. (Rest of panels) Sample time signals for a single actuator in: opposition control (blue), DRL for drag reduction (orange), DRL for Q events reduction (green), DRL for streak reduction (red), and DRL for SHAP reduction (purple).
An additional advantage of our XDL approach is observed when analyzing control dynamics and energy requirements. The power input of the control is defined as:
$w_{\text{in}}=\frac{1}{2}||v_{\text{wall}}||^{3}$ (5)
for each agent. Fig. 4 shows sample actuation signals from the different control strategies, along with their associated power input requirements. Direct drag reduction control and streak-based control frequently oscillate between the minimum and maximum allowed control values in a “bang-bang” pattern, resulting in high power input. The SHAP-based control operates more efficiently, rarely reaching maximum or minimum control values and exhibiting fewer extreme actuation patterns. This translates to considerable input power savings: the SHAP-based approach requires only half the power input compared with the DD control while delivering better drag reduction. The Q-event-based control, while even more energy-efficient, yields significantly lower drag reduction than the SHAP-based approach, as discussed above.
The net energy saving is defined as:
$S=\frac{c_{f,\text{uncontrolled}}-(c_{f}+w_{\text{in}})}{c_{f,\text{uncontrolled}}},$ (6)
[…]
Conclusions
This study demonstrates a novel approach to turbulence control by integrating explainable deep learning (XDL) with deep reinforcement learning (DRL) using the Shapley additive explanations (SHAP values). The SHAP values identify the most important coherent structures in the flow and they are capitalized on to discover control strategies for wall-bounded turbulence. By training DRL models in a minimal flow configuration with a single self-sustaining process, we successfully developed control strategies which can be deployed to much larger channels with multiple interacting SSPs.
Our analysis reveals several critical insights. While the DRL approach targeting Q events excels at reducing the classical Q events and streaks, it falls short in reducing those structures truly relevant for drag reduction. In contrast, the SHAP-based approach, relying on the SHAP values, yields a more effective control strategy, achieving $33.3\%$ drag reduction. This represents a $5\%$ improvement over the DRL approach directly targeting drag reduction, a $21.2\%$ improvement over Q-event-based control, and a $57\%$ improvement over the streak-based control. Furthermore, the SHAP-based control only requires half of the power input compared with the DRL approach targeting drag reduction.
It is also interesting to note that in the SHAP-based approach all input information comes from a single plane parallel to the wall and that all training is performed with just a single SSP, and consequently, the amount of experiences observed in training is limited. This approach demonstrates the potential of XDL in identifying critical flow mechanisms, where the proposed method is capable of capturing the causal relationships between flow structures, rather than targeting individual coherent structures in isolation.
The framework introduced in this work is a promising entry point for the powerful and unique combination of DRL and XDL, where there is still a wide range of possible improvements such as: usage of a convolutional neural network in the DRL policy (instead of a multilayer perceptron), training using volumetric data instead of a single plane, or transfer learning from the small training domain to the large channel one. The use of XDL for DRL can have a great impact in a wide range of applications as it can be applied directly on experimental data [24], avoiding the ever growing requirements for computational power and high-resolution simulation. The SHAP reward introduced in this work provides new insights into the most critical mechanisms of turbulence and provides new paths to tame and harness the power of turbulent flows in many industrial (e.g. aerodynamics and food production), biological (e.g. active matter) and medical applications (e.g. personalised medicine).
Methods
Computational set-up
As introduced in the main text, we consider a turbulent open-channel flow of height $h$, driven by a time-varying pressure gradient that maintains a constant mass flux. The spatial coordinates are $x$, $y$, and $z$ in the streamwise, wall-normal, and spanwise directions, respectively. We solve the Navier–Stokes equations, which are made non-dimensional with the open-channel height $h$ and the laminar centerline velocity $U_{cl}$, and read:
$\partial_{t}\mathbf{u}+(\mathbf{u}\cdot\nabla)\mathbf{u}$ $=-\nabla p+\frac{1}{\textit{Re}}\Delta\mathbf{u},$ (7)
$\nabla\cdot\mathbf{u}$ $=0.$ (8)
The flow field is entirely described by the velocity vector $\mathbf{u}(x,y,z,t)=(u,v,w)$. Direct numerical simulations (DNS) for training are carried out in a computational domain of dimensions $[L_{x},L_{y},L_{z}]=[2.67,1,0.8]$, i.e. the small channel configuration (SCC), and the evaluation of the resulting policies is performed in a box of size $[L_{x},L_{y},L_{z}]=[2\pi,1,\pi]$, i.e the large channel configuration (LCC). We consider Dirichlet boundary conditions at the lower wall and symmetry boundary conditions at the upper boundary:
$u(y=0)=w(y=0)=0,\ v(y=0)=v_{bc},$ (9)
$\partial_{y}u(y=1)=\partial_{y}w(y=1)=0,\ v(y=1)=0.$ (10)
The uncontrolled turbulent flow is characterized by the friction Reynolds number, which is set to $Re_{\tau}=180$.
To perform our DNS we use the generic partial differential equation (PDE) solver Dedalus [31]. This solver uses a pseudo-spectral method, where solutions are expanded into Fourier series in the homogeneous ($x$ and $z$) directions and into Chebyshev polynomials into the inhomogeneous (wall-normal, $y$) direction. We expand solutions of the SCC into $[N_{x},N_{y},N_{z}]=[16,64,16]$ modes and we timestep the simulations using a third-order four-stage diagonally implicit Runge–Kutta and explicit Runge–Kutta (DIRK+ERK) scheme [32] with a constant time step $\Delta t^{+}\approx 0.039$ ($\Delta t=0.005$). These computational parameters are common in the literature for the moderate Reynolds number considered in this study [22]. The LCC simulations are performed expanding the solutions into $[N_{x},N_{y},N_{z}]=[64,64,32]$ modes using the same time-stepping scheme and time step as in the SCC case. These configurations (SCC and LCC) are used to perform long simulations ($t^{+}>20,000$) which form the dataset for training (and testing) required for the U-net models and the DRL framework.
Reinforcement learning configuration and training
As briefly described in the introduction, a DRL set-up consists of two main elements that interact with each other: the environment, i.e. the system where we input actions, obtain rewards, and observe states, and the agent, i.e. the controller in charge of deciding the actions to take based on inputs from the environment. In the present work, as shown in Fig. 1, the environment corresponds to a numerical simulation performed using the Dedalus codebase [31], which accounts for the main computational cost in the DRL set-up. To address this, the Dedalus code is parallelized over several processors using the message-passing interface (MPI).
As mentioned, the DRL setup requires the choice of an agent, which can be classified as model-based or model-free. Model-based agents are based on a model of the dynamics of the environment, while model-free methods optimize the agent in a trial-an-error process carrying out a number of episodes. In the present work we choose a model-free agent. Such agents can be further split into policy-gradient, value-function, and a combination of policy-gradient and value-function algorithms dubbed actor-critic algorithms. A policy-gradient algorithm is based on the parametrization of the agent and its optimization to maximize cumulative rewards. Value-function algorithms, in contrast, attempt to estimate the cumulative reward given a state (state value function $V^{\pi}(s)$) or a state-action couple (action value function $Q^{\pi}(s,a)$), where:
$V^{\pi}(s)$ $=\mathbb{E}_{\pi}\left[r|s\right],$ (11)
$Q^{\pi}(s,a)$ $=\mathbb{E}_{\pi}\left[r|s,a\right],$ (12)
with $\mathbb{E}$ denoting the expectation. Moreover, model-free agents can be classified as on-policy or off-policy. In on-policy methods, the optimization and policy update are based on the learning generated by the current agent. Off-policy algorithms use a replay buffer to store trajectories generated with previous policies, and these are used to perform the policy update. In the present work we have chosen the model-free, off-policy TD3 algorithm [33] as implemented in the Stable Baselines 3 library [34]. The combination of the Dedalus codebase with Stable Baselines 3 is particularly convenient since both of them are Python based. Note that this is a robust state-of-the-art DRL algorithm suitable for continuous control problems [33]. The reinforcement learning side of the set-up is much less computationally demanding and can be efficiently trained on a single CPU.
In the physical configuration for the SCC described above, we defined the same number of controllers (agents) as grid points (modes) in the periodic directions, such that each agent sees only the state $(u,v)$ located at $y^{+}=15$ right above. The spatial average of the state is removed at each step, i.e. whenever information is communicated to the agent to avoid biasing the observations when evolving the flow between the controlled and uncontrolled cases (note that this is only required when information is passed to the agent). Consequently, each actuator is defined as a multi-agent reinforcement-learning pseudo-environment, resulting in $N_{x}\times N_{z}=256$ independent trajectories that the agent uses during the optimization step. The actions taken by the agents correspond to blowing and suction, i.e. positive/negative values of the vertical velocity at the wall, and they are limited to the range $[-u_{\tau},u_{\tau}]$. In order to ensure that no net mass is introduced by the controllers, we enforce a zero-mean action over all agents. This is done by removing the spatial average over all actuators before the control is applied in the DNS. The control interacts with the DNS every $\Delta t^{+}\approx 0.54$ ($\Delta t=0.07)$ and each episode consists of $3,000$ interactions, in agreement with prior studies [22]. Between consecutive agent interactions, the instantaneous values of the reward are calculated and stored at each DNS step. These stored values are then averaged to produce a single reward value.
We train each model using 6 different initial conditions in the SCC. At the start of each episode a random number generator chooses which initial condition is used for that particular trajectory. Since the TD3 algorithm uses a deterministic policy in training, we introduce noise with amplitude $0.1u_{\tau}$ to favor exploration. During training in the SCC, relaminarization of the flow can be observed and in such cases we terminate the episodes once the flow is deemed to approach the laminar state sufficiently, i.e. once the drag reduction is over $55\%$. This is done to avoid providing the model with very large rewards for actions which no longer affect the flow on its path to relaminarization. The model is evaluated every 5 episodes on an additional initial condition not seen in training. The best performing model on the evaluated initial condition is then stored. Albeit several evaluation episodes could be considered, this process becomes expensive and a single evaluation initial condition proved to be sufficient to identify effective DRL models. Moreover, we perform several training runs to explore a wider range of control strategies. The Supplementary material shows the reward metrics during sample training runs for the different rewards considered. It can be observed that typically the model improves quickly at the start of the process and then discovers many similarly effective policies during training. Throughout the training, the model passes through parameter regions where better policies are found, but also where it unlearns. Each 100 episodes corresponds to roughly 96 hours of computing using four AMD Ryzen 9 7950X CPUs.
The best policy for each of the four DRL rewards in the SCC is tested in 50 initial conditions randomly sampled from our database of $30,000$ instantaneous fields and unseen during training as reported in the Supplemetary material. The results show that all DRL policies outperform opposition control and for $t^{+}>500$ they achieve drag reductions of: $36.5\%$ for the DD reduction reward, $36.2\%$ for the Q-event-based reward, $34.4\%$ for the streak-based reward, and $37.4\%$ for the SHAP-based reward. It can also be observed that in this small channel, the SHAP-based reward produces the best performing policy in terms of reducing the Q events and streaks.
Calculation of the XDL-based reward
Defining a criterion for calculating the reward requires determining the regions of the flow with a higher impact on its evolution. In order to calculate these high-importance regions, an explainable-deep-learning (XDL) approach is employed. This approach is based on three steps: i) predict a future state of the flow through a deep-learning model, ii) calculate the importance of each grid point in the prediction of flow in a future time and iii) train a model that predicts the field of importance (SHAP values) from the original velocity fluctuation field.
For the first stage, a U-net architecture [35], $f_{u}$, is employed to predict the velocity fluctuation field in a future time step ($\mathbf{u}^{p}_{t+\Delta t}$) with $\Delta t^{+}=\Delta t\,u_{\tau}^{2}/\nu\approx 5$, from the true velocity field at the present time ($\mathbf{u}_{t}$): $\mathbf{u}^{p}_{t+1}=f_{u}\left(\textbf{u}_{t}\right)$. Both input and output fields have a size of $16\times 64\times 16$ grid points in the streamwise, wall-normal and spanwise directions, respectively. In addition, a periodic padding of $3$ is added in the streamwise and spanwise directions. The U-net has approximately $1.5$ million parameters, distributed on $4$ levels in the U-net, with $16$, $32$, $64$, and $128$ filters and a kernel of size $3\times 3$ from the first to the last level. Each level is composed of two blocks (3D convolutional layer+batch normalization+activation function) and one block in the decoder. The encoder levels are connected by average poolings with size $2$, while in the decoder, the levels are connected with transposed convolutions with $16$, $32$, and $64$ filters and kernel of size $3\times 3$ from top to bottom level. Each level of the encoder and the decoder are connected by concatenating the output of the encoder level with the output of the transposed convolution. The U-net is trained on a database of $30,000$ instantaneous flow fields, using $20\%$ for validation, until the prediction error is approximately $1\%$.
Once the model for predicting the velocity is trained, the importance of each grid point for the accuracy of the prediction is evaluated. For this reason, the model $f_{u}$ is modified in order to calculate the mean-squared error of the predictions:
${\rm MSE}_{t+\Delta t}=f_{\rm MSE}\left(\mathbf{u}_{t}\right)=\frac{1}{N_{x}N_{y}N_{z}}\sum_{i_{x}=1}^{N_{x}}\sum_{i_{y}=1}^{N_{y}}\sum_{i_{z}=1}^{N_{z}}\left(\mathbf{u}_{t+\Delta t}-f_{u}\left(\mathbf{u}_{t}\right)\right)^{2}\rm{.}$ (13)
Then, the importance of each grid point is calculated using additive-feature-attribution methods [25], which substitute the deep-learning model $f_{\rm MSE}$ by a surrogate linear model $g_{\rm MSE}$ defined as:
${\rm MSE}_{t+\Delta t}=f_{\rm MSE}\left(\mathbf{u}_{t}\right)\approx g_{\rm MSE}\left(z_{ji}\right)=\phi_{0}+\sum_{j\in(u,v,w)}\sum_{i=0}^{N}\phi_{ji}z_{ji}\rm{,}$ (14)
where the importance, or Shapley additive explanations (SHAP) values, of each single grid point $i$ for the component $j$ of the velocity is defined by $\phi_{ji}$ and $N$ is the total number of grid points in a single snapshot. The parameter $z_{ji}$ is a boolean value which is 00 or $1$ in case of removing or including the information of the velocity component, $j$, in this specific grid point, $i$. The expected mean-squared error of the prediction when all the grid points are removed is defined by $\phi_{0}$. Note that the linear model presented in equation (14) is a local approximation of the MSE, and thus, it is updated for every snapshot. However, the computational cost of these values increases exponentially with the number of parameters, i.e. as $2^{N}$ [36]. To reduce the computational cost of the calculations, the prior knowledge of the mathematical definition of the architecture and the neurons can be exploited [37]. In the present work, the gradient-SHAP algorithm, an additive-feature-attribution method for differentiable models based on the expected-gradients method [38], is used to simplify the SHAP-value calculation as follows:
$\phi_{ji}\left(u_{t}\right)=\mathbb{E}_{u_{\rm{ref}},\alpha\sim U\left(0,1\right)}\left[\left(u_{t_{ji}}-u_{t_{ji}}\right)\frac{\partial f_{\rm MSE}\left(u_{\rm{ref}}+\alpha\left(u_{t}-u_{\rm{ref}}\right)\right)}{\partial u_{t_{ji}}}\right]\text{,}$ (15)
where $u_{\rm{ref}}$ is the reference velocity field. For this analysis, the instantaneous spatially-averaged mean flow is used as a reference as the instantaneous mean value of the velocity is non-informative for the evaluation of the error of the model.
After the SHAP values of the $30,000$ instantaneous flow fields are calculated, a new U-net is trained to predict the SHAP values from the present velocity fluctuation field: $\bm{\phi}=f_{\phi}\left(\mathbf{u}_{t}\right)$. The architecture of this model is identical to that of the previous U-net ($f_{u}$), using in this case the instantaneous velocity field as input and the SHAP-value field as output. A database comprising the previous $30,000$ instantaneous flow fields is used to train the model, reserving $20\%$ for validation, until the prediction error is lower than $0.5\%$. Finally, the SHAP-value field $\bm{\phi}$ is predicted in order to compute the reward of the DRL in the simulation loop with a low computational effort compared to that of explicitly computing the SHAP values, making this step over 200 times faster.
[…]
Acknowledgements
The deep-learning-model training was enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Dardel (PDC) and Alvis (CS3E). M.B. and R.V. acknowledge partial financial support from Digital Futures. A.C. and R.V. acknowledge financial support from ERC grant no. ‘2021-CoG-101043998, DEEPCONTROL. Views and opinions expressed are those of the authors only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
[…]
Author contribution Beneitez, M.: Methodology, Software, Validation, Investigation, Writing - Original Draft, Visualization. Cremades, A.: Methodology, Software, Writing - Original Draft. Guastoni, L.: Methodology, Software, Writing - Original Draft. Vinuesa, R.: Conceptualization, Project definition, Methodology, Resources, Writing - Original Draft, Supervision, Project administration, Funding acquisition.